OpenStudy (anonymous):

How do I write an equation of a line, in slope-intercept form, that has slope 3/2 and passes through (0,-7)

OpenStudy (hoblos):

the general formula is y=mx+b where m is the slope and b is the y-intercept

OpenStudy (anonymous):

y=mx+c so y=3/2x -7

OpenStudy (hoblos):

so here y= (3/2)x -7

OpenStudy (anonymous):

so that would be the equation that would be used?

OpenStudy (hoblos):

yes thats it... y= (3/2)x -7

OpenStudy (anonymous):

Thanks

OpenStudy (hoblos):

yw :)